A floor is 5 m long and 4 m wide. A square carpet of sides 3 m is laid on the floor. Find the area of the floor that is not carpeted.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Area of floor = l × b = 5 × 4
= 20 m^2
Area of square carpet = 3 × 3
= 9 m^2
Area of floor that is not carpeted = 20–9
= 11 m^2
∴Area of the floor that is not carpeted is 11 m^2

Thanks